OpenStudy (anonymous):

An apartment complex contains 230 apartments each having 1, 2 or 3 bedrooms. The number of 2-bedroom apartments is 10 more than 3 times the number of 3-bedroom apartments. The number of 1-bedroom apartmenst is twice the number of 2-bedroom apartments. How many apartments of each kind are in the complex? Please answer with procedure

jimthompson5910 (jim_thompson5910):

Let x = number of 1-bedroom apartments y = number of 2-bedroom apartments z = number of 3-bedroom apartments

jimthompson5910 (jim_thompson5910):

`The number of 2-bedroom apartments is 10 more than 3 times the number of 3-bedroom apartments` gives the equation y = 3z + 10

jimthompson5910 (jim_thompson5910):

`The number of 1-bedroom apartmenst is twice the number of 2-bedroom apartments` means x = 2y

jimthompson5910 (jim_thompson5910):

`An apartment complex contains 230 apartments each having 1, 2 or 3 bedrooms` so x+y+z = 230

jimthompson5910 (jim_thompson5910):

you have this system of equations \[\Large \begin{cases}y = 3z + 10\\x = 2y\\x+y+z = 230\end{cases}\]

jimthompson5910 (jim_thompson5910):

Do you see how to solve?

OpenStudy (anonymous):

Honestly i dont...

OpenStudy (anonymous):

Its the only one i cant seem to get

jimthompson5910 (jim_thompson5910):

since y = 3z+10, we can replace the y in x = 2y with 3z+10 x = 2y x = 2*(3z+10) ... replace y with 3z+10 x = 6z+20

jimthompson5910 (jim_thompson5910):

x+y+z = 230 6z+20+y+z = 230 ... replace x with 6z+20 6z+20+3z+10+z = 230 ... replace y with 3z+10

jimthompson5910 (jim_thompson5910):

now solve `6z+20+3z+10+z = 230` for z

OpenStudy (anonymous):

omg thanks brother You helped me out a lot today

jimthompson5910 (jim_thompson5910):

no problem